# Neural Uncertainty Principle: A Unified View of Adversarial Fragility and LLM Hallucination

### 26 Mar 2026 v2

@@ VI-A A conjugate limit underlying boundary behaviors
Interpreting the operator moments as second-order geometry of the loss-weighted $(x_{u},p_{u})$ cloud (Sec. III-D) turns the inequality into a geometric statement: the effective area $\sqrt{\det\Sigma(u)}=\Delta\hat{m}_{u}^{\star}\,\Delta\hat{p}_{u}$ is lower-bounded (Eq. (26)). When the system operates near this limit, improvements in one axis necessarily incur costs in the conjugate axis. This offers a principled explanation for why purely optimizing one desideratum (e.g., squeezing boundary errors to raise accuracy) can systematically amplify another failure mode (e.g., sensitivity exploited by adversarial perturbations).
@@ VI-B Unified interpretation across vision and language
In discriminative vision, training can push hard samples toward a boundary-stress configuration: the ambiguity thickness becomes small while sensitivity dispersion becomes large, with strong input–gradient coupling concentrating on a stressed tail (Exp. 1–2; Proposition 12). In LLM prompting under a prefill-only constraint, the dominant pathology is instead under-conditioning: unusually weak prompt–gradient coupling indicates that the prompt does not effectively constrain loss-sensitive directions, leaving high slack and enabling prior-driven drift (Exp. 5–6; Proposition 4). The same coupling channel therefore provides a common language for boundary anomalies across perception and generation.
@@ VI-C From a bound to practical observables and control channels
A second theoretical contribution is the bridge from the operator inequality to computable quantities. Under the loss-induced state, the symmetrized operator covariance reduces to a scalar covariance of real gradients (Theorem III.2), and random directional probing links projected correlation to cosine (Lemma III.2). This justifies the CC-Probe as a single-backward proxy for the coupling channel that modulates proximity to the uncertainty limit. Once coupling is made observable, it also becomes controllable: ConjMask directly suppresses dominant per-component couplings $|x_{i}p_{i}|$ and yields strong robustness under PGD-20 and APGD-CE without adversarial training (Exp. 3; Proposition 3). The CE–DLR gap further indicates that robustness is multi-channel; LogitReg complements ConjMask by stabilizing score-space geometry and broadening robustness under APGD-DLR (Exp. 4).
@@ VI-D Limitations and future directions
Our analysis is most informative when there exists a nontrivial boundary-relevant subset that the loss-phase construction can emphasize. In the language setting, the usefulness of CC-Probe depends on a nontrivial match between benchmark difficulty and model capability: if the benchmark is too easy hallucinations are too rare, if it is too hard failures become nearly ubiquitous, and the probe is therefore most informative on prompts near the model’s effective uncertainty boundary.
In near “pure fitting” regimes—e.g., when the loss is essentially small on the relevant support, or the input–output mapping behaves close to bijective with little effective compression [40]—the weighting $w_{c}(x)\propto\mathcal{L}_{c}(x)^{2}$ provides little contrast and distinct loss-induced states $\psi_{c}$ are not meaningfully separated; consequently, the notion of a stressed boundary layer becomes ill-posed and the CC-Probe inclines to lose discriminative power. In addition, the probe is loss-dependent (CE in vision; shifted prompt NLL in language), and Exp. 3 indicates robustness can differ across attack losses (CE vs. DLR), motivating multi-loss evaluation (e.g., APGD-CE and APGD-DLR) and defenses that cover multiple geometry channels. Finally, CC-Probe is a proxy rather than a direct estimator of $\rho_{c}(u)$ or slack $S_{c}(u)$, and prefill-only LLM use still requires one backward pass on prompt embeddings. Future work includes designing alternative state constructions when $\mathcal{L}_{c}$ is near-degenerate, developing stable estimators of direction-dependent slack, and extending coupling-based monitoring/control to multimodal and preference-optimized settings.
Crucially, the NUP framework also serves as a predictive warning for future LLM alignment. Currently, LLM hallucinations are predominantly driven by a high-slack (under-conditioned) regime, as detected by the CC-Probe. However, if future training paradigms aggressively reduce this slack to suppress hallucinations—forcing prompts into tight, deterministic couplings with specific continuations—the NUP inequality $\Delta\hat{m}_{u}^{\star}\,\Delta\hat{p}_{u}\geq\tfrac{1}{2}$ dictates that the system could inevitably shift toward the opposite extreme. In doing so, we risk inducing the same boundary-stress regime currently plaguing vision models, which would manifest in LLMs as extreme adversarial fragility (e.g., imperceptible prompt perturbations or synonymous substitutions triggering catastrophic output shifts). Ultimately, the NUP demonstrates that we cannot entirely eliminate boundary uncertainty; we can only trade one form of fragility for another. Sustainable model design might therefore navigate this fundamental trade-off, seeking an intermediate “Goldilocks” zone rather than minimizing a single failure axis.
@@ VII Conclusion
In this study we have introduced the Neural Uncertainty Principle (NUP), a fundamental geometric constraint showing that input projections and their loss gradients are conjugate observables under a loss-induced state. NUP unifies adversarial vulnerability in vision and hallucination in LLMs as opposite extremes on the same conjugate plane—saturation versus slack. Based on this principle, we derive a practical single-backward probing method, CC-Probe, which enables the diagnosis of cross-modal boundary anomalies and supports targeted interventions: ConjMask and LogitReg for enhancing robustness in vision tasks, and prefill-stage risk scoring for hallucination detection and prompt selection in language models. By reframing two isolated reliability challenges as a single conjugate trade-off, the NUP provides a unified perspective for advancing robust and faithful AI.

### 27 Mar 2026 v3

@@ VI-A A conjugate limit underlying boundary behaviors
Interpreting the operator moments as second-order geometry of the loss-weighted $(x_{u},p_{u})$ cloud (Sec. III-D) turns the inequality into a geometric statement: the effective area $\sqrt{\det\Sigma(u)}=\Delta\hat{m}_{u}^{\star}\,\Delta\hat{p}_{u}$ is lower-bounded (Eq. (26)). When the system operates near this limit, improvements in one axis necessarily incur costs in the conjugate axis. This offers a principled way to interpret why optimizing one desideratum (e.g., squeezing boundary errors to raise accuracy) can systematically amplify another failure mode (e.g., sensitivity exploited by adversarial perturbations).
@@ VI-B Unified interpretation across vision and language
In discriminative vision, training can push hard samples toward a boundary-stress configuration: the ambiguity thickness becomes small while sensitivity dispersion becomes large, with strong input–gradient coupling concentrating on a stressed tail (Exp. 1–2; Proposition 12). In LLM prompting under a prefill-only constraint, the dominant pathology is instead under-conditioning: unusually weak prompt–gradient coupling indicates that the prompt does not effectively constrain loss-sensitive directions, leaving high slack and enabling prior-driven drift (Exp. 5–6; Proposition 4).
@@ VI-C From a bound to practical observables and control channels
A second theoretical contribution is the bridge from the operator inequality to computable quantities. Under the loss-induced state, the symmetrized operator covariance reduces to a scalar covariance of real gradients (Theorem III.2), and random directional probing links projected correlation to cosine (Lemma III.2). This justifies the $C$-Probe as a single-backward proxy for the coupling channel that modulates proximity to the uncertainty limit. Once coupling is made observable, it can also be used as a practical control target: ConjMask directly suppresses dominant per-component couplings $|x_{i}p_{i}|$ and yields strong robustness under PGD-20 and APGD-CE without adversarial training (Exp. 3; Proposition 3). The CE–DLR gap further indicates that robustness is multi-channel; LogitReg complements ConjMask by stabilizing score-space geometry and broadening robustness under APGD-DLR (Exp. 4).
@@ VI-D Limitations and future directions
Our analysis is most informative when there exists a nontrivial boundary-relevant subset that the loss-phase construction can emphasize. In near “pure fitting” regimes—e.g., when the loss is essentially small on the relevant support, or the input–output mapping behaves close to bijective with little effective compression [40]—the weighting $w_{c}(x)\propto\mathcal{L}_{c}(x)^{2}$ provides little contrast and distinct loss-induced states $\psi_{c}$ are not meaningfully separated; consequently, the notion of a stressed boundary layer becomes ill-posed and the $C$-Probe inclines to lose discriminative power. In addition, the probe is loss-dependent (CE in vision; shifted prompt NLL in language), and Exp. 3 indicates robustness can differ across attack losses (CE vs. DLR), motivating multi-loss evaluation (e.g., APGD-CE and APGD-DLR) and defenses that cover multiple geometry channels. Finally, $C$-Probe is a proxy rather than a direct estimator of $\rho_{c}(u)$ or slack $S_{c}(u)$, and prefill-only LLM use still requires one backward pass on prompt embeddings.
The NUP framework also suggests a cautionary perspective for future LLM alignment. Currently, LLM hallucinations are predominantly driven by a high-slack (under-conditioned) regime, as detected by the $C$-Probe. However, if future training paradigms aggressively reduce this slack to suppress hallucinations—forcing prompts into tight, deterministic couplings with specific continuations—the NUP inequality $\Delta\hat{m}_{u}^{\star}\,\Delta\hat{p}_{u}\geq\tfrac{1}{2}$ dictates that the system may be pushed toward the opposite extreme. In doing so, we risk inducing the same boundary-stress regime currently plaguing vision models, which would manifest in LLMs as extreme adversarial fragility (e.g., imperceptible prompt perturbations or synonymous substitutions triggering catastrophic output shifts). Ultimately, the NUP suggests that boundary uncertainty may not be fully eliminable within the present formulation; in practice, reducing one form of fragility may increase pressure on another. Sustainable model design might therefore navigate this fundamental trade-off, seeking an intermediate “Goldilocks” zone rather than minimizing a single failure axis.
@@ VII Conclusion
In this study, we introduced the Neural Uncertainty Principle (NUP), a geometric constraint-based perspective in which input projections and their loss gradients behave as conjugate observables under a loss-induced state. NUP unifies adversarial vulnerability in vision and hallucination in LLMs as opposite extremes on the same conjugate plane—saturation versus slack. Based on this principle, we derive a practical single-backward probing method, $C$-Probe, which enables the diagnosis of cross-modal boundary anomalies and supports targeted interventions: ConjMask and LogitReg for enhancing robustness in vision tasks, and prefill-stage risk scoring for hallucination detection and prompt selection in language models. By reframing two isolated reliability challenges as a single conjugate trade-off, the NUP provides a unified perspective for advancing robust and faithful AI.